Operations Report – December 21st

Crew 236 Operations Report 21Dec2021

SOL: 2

Name of person filing report: Pavithra Ravi

Non-nominal systems:

Notes on non-nominal systems:

Generator:

Hours run: 13 hours 48 minutes

From what time last night: 18:07

To what time this morning: 07:55

List any additional daytime hours when the generator was run: None

Solar— SOC % (Before generator is run at night): Information provided by Outpost

Station Propane Reading – % Information provided by Outpost

Ethanol Free Gasoline: Not in use

Water (loft tank): 29 gallons

Water Meter: 1539449 units

Water (static tank): approx. 415 gallons

Static to Loft Pump used: No

Water in GreenHab: approx. 232 gallons

Water in ScienceDome: N/A

Toilet tank emptied: No

Deimos rover used: No

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Spirit rover used: No

Hours (before EVA 1): N/A

Beginning charge (Before EVA 1): N/A

Ending charge (On return from EVA 1, before recharging): N/A

Hours (before EVA 2): N/A

Beginning charge (Before EVA 2): N/A

Ending charge (On return from EVA 2, before recharging): N/A

Currently charging: Yes

Opportunity rover used: No

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: Yes

Curiosity rover used: No

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: Yes

Notes on rovers: None

ATV’s Used (Honda, 300, 350.1, 350.2, 350.3): None

Reason for use: N/A

Oil Added? N/A

ATV Fuel Used: N/A

# Hours the ATVs were used today: N/A

Notes on ATVs: N/A

HabCar used and why, where? No

CrewCar used and why, where? No

General notes and comments:

Today the toilet leak worsened, requiring frequent dumping out of the
collected water. This has now been dealt with by keeping the water
valve closed, except for when flushing is required.

The exposed power outlets above the stove in the hab have been covered
with gorilla tape as a safety precaution.

The gaps surrounding the window above the couch in the hab have also
been blocked out with gorilla tape to minimize influx of (lethal)
martian atmosphere into our living area.

The temperature on the upper level of the hab has thankfully been
stabilized at around 70 F. The random bursts of cool air which we
received yesterday seem to have stopped (the trick was keeping the
thermostat at 70 instead of 75).

Summary of internet: Nominal

Summary of suits and radios: Nominal

Summary of Hab operations: Nominal.

Summary of GreenHab operations: Nominal

Summary of ScienceDome operations: Nominal

Summary of RAM operations: Nominal

Summary of any observatory issues: In correspondence with the
astronomy support team.

Summary of health and safety issues: Nominal

Questions, concerns and requests to Mission Support: None